The Belgian food group Delhaize just paid a dividend of €1.08. This is paid in semiannual instalments. The firm has a policy to pay out 75 per cent of its annual dividend after six months and the remaining amount at the end of the year. If the annual discount rate on Delhaize shares is 8 per cent and the dividend is expected to grow at 3 per cent per year, what is the current share price of Delhaize? [Hint, treat these are two growing perpetuity streams and it paid out €0.81 six months ago]
Answer #1

Present Value of growing Perpetuity = D1 / (r-g)

where, D = Dividend at period 1, r = discount rate and g = growth rate

So, PV = 0.81 / (0.08-0.03) = Euro 16.2

Current Share Price of Delhaize is Euro 16.2